# Self-supervised learning of dense hierarchical representations for medical image segmentation

### v1

@@ 1 Introduction
Our major contributions are as following: (1) We propose a self-supervised framework for learning dense hierarchical representations that effectively mitigates the intrinsic bias of the FPN-based approach towards prioritizing global context features over the fine-grained features. (2) We demonstrate that our strategy is particularly beneficial for downstream segmentation tasks with limited annotated data, and show that our pre-trained model significantly outperforms the baseline counterpart in linear evaluation setting. (3) We publicly release the pre-trained models for both CT and MRI modalities. For the best of our knowledge we are the first to release a model pre-trained on a large collection of MRI images.
@@ 3.3 Results
For the experiments with the AMOS MRI dataset we employ a 5-fold cross validation, and a 3 fold cross-validation for the experiments with the BTCV CT dataset. Performance of all models is evaluated using the Dice score.
aug $L_{c}$ $L_{r}$ arch Sp Kid Gb Es Li St Aor IVC Pa AG Avg
@@
First, to assess the impact of individual components on the final performance of our method we conduct ablation experiments, employing linear evaluation settings and MRI data (Table 1). We start from the baseline vox2vec [11] model that contains only contrastive loss function $L_{c}$ and evaluate different combinations of the proposed modifications: local region augmentations, restorative loss function $L_{r}$ and architectural changes. Remarkably, addition of each of these individual components consistently lead to performance improvement. The results confirm the importance of each modification in mitigating the inherent imbalance of hierarchical contrastive learning.
@@
Table 2: Results of linear evaluation for CT data. We employed 2500 images to pre-train our models and compared their performance with publicly available vox2vec model which was pre-trained using 6550 images.
@@
method data Sp Kid Gb Es Li St Aor IVC Pa AG Avg
from scratch 42 89.3 90.7 59.6 61.0 93.7 81.1 89.5 84.4 76.9 54.4 77.1 $\pm$ 5.4
fine-tuning 42 90.8 91.7 67.0 62.3 94.5 81.9 89.6 84.9 79.1 56.2 78.8 $\pm$ 4.3
from scratch 8 81.3 80.6 34.5 47.0 90.7 65.7 80.8 69.3 57.5 35.6 63.3 $\pm$ 6.7
fine-tuning 8 86.8 85.6 50.8 46.0 93.0 66.9 84.6 76.7 67.8 44.3 69.3 $\pm$ 4.7
from scratch 4 66.2 67.7 18.7 35.0 85.1 40.7 72.1 54.9 39.2 16.6 48.4 $\pm$ 10.1
fine-tuning 4 81.2 78.6 20.8 41.0 90.9 52.7 78.3 63.6 51.1 27.5 57.7 $\pm$ 7.8
Table 3: Results of fine-tuning for MRI data. We conducted experiments using 42, 8, and 4 images and compared the performance of the fine-tuned models to a models trained from scratch.

### v2

@@ 1 Introduction
Our major contributions are as following: (1) We propose a self-supervised framework for learning dense hierarchical representations that effectively mitigates the intrinsic bias of the FPN-based approach towards prioritizing global context features over the fine-grained features. (2) We demonstrate that our strategy is particularly beneficial for downstream segmentation tasks with limited annotated data, and show that our pre-trained model significantly outperforms the baseline counterpart in linear evaluation setting. (3) We publicly release the pre-trained models for both CT and MRI modalities.
@@ 3.3 Results
For the experiments with the AMOS MRI dataset we employ a 5-fold cross validation, and a 3 fold cross-validation for the experiments with the BTCV CT dataset. Performance of all models is evaluated using the Dice score. For all experiments we report mean results of cross-validation for the abdominal organs: spleen $Sp$, kidneys $Kid$, gallbladder $Gb$, esophagus $Es$, liver $Li$, stomach $St$, aorta $Aor$, inferior vena cava $IVC$, pancreas $Pa$ and adrenal glands $AG$. We also report the mean and standard deviation of the overall Dice score across folds.
aug $L_{c}$ $L_{r}$ arch Sp Kid Gb Es Li St Aor IVC Pa AG Overall
@@
First, to assess the impact of individual components on the final performance of our method we conduct ablation experiments, employing linear evaluation settings and MRI data (Table 1). We start from the baseline vox2vec [11] model that contains only contrastive loss function $L_{c}$ and evaluate different combinations of the proposed modifications: local region augmentations $aug$, restorative loss function $L_{r}$ and architectural changes $arch$. Remarkably, addition of each of these individual components consistently lead to performance improvement. The results confirm the importance of each modification in mitigating the inherent imbalance of hierarchical contrastive learning.
@@
Table 2: Results of linear evaluation on CT data. We employed 2500 images to pre-train model with our method and vox2vec method, and compared their performance with publicly available vox2vec model which was pre-trained using 6550 images.
@@
method Sp Kid Gb Es Li St Aor IVC Pa AG Overall
42 training images
from scratch 89.3 90.7 59.6 61.0 93.7 81.1 89.5 84.4 76.9 54.4 77.1 $\pm$ 5.4
fine-tuning 90.8 91.7 67.0 62.3 94.5 81.9 89.6 84.9 79.1 56.2 78.8 $\pm$ 4.3
8 training images
from scratch 81.3 80.6 34.5 47.0 90.7 65.7 80.8 69.3 57.5 35.6 63.3 $\pm$ 6.7
fine-tuning 86.8 85.6 50.8 46.0 93.0 66.9 84.6 76.7 67.8 44.3 69.3 $\pm$ 4.7
4 training images
from scratch 66.2 67.7 18.7 35.0 85.1 40.7 72.1 54.9 39.2 16.6 48.4 $\pm$ 10.1
fine-tuning 81.2 78.6 20.8 41.0 90.9 52.7 78.3 63.6 51.1 27.5 57.7 $\pm$ 7.8
Table 3: Comparison of fine-tuning with training from scratch on MRI data of various sizes: 42, 8, and 4 training images.